Show work

1. Convert 0.025 kg/L to cg/dL (6 pts)

2. The molecular formula of ethanol is C2H6O. The density of ethanol is 0.997 g/mL. Calculate the volume (in mm3) of 5.7 grams of ethanol. (8 pts)

3. A bubble of helium gas has a volume of 0.650 mL near the bottom of a large aquarium where the pressure is 1.54 atm and the temperature is 12°C. Determine the bubble’s volume upon rising near the top where the pressure is 1.01 atm and 16°C. Assume that the number of moles of helium remains constant and that the helium is an ideal gas. (8 pts)

4. When strontium nitrate is heated in fireworks, it gives off a red light that has a wavelength of 650 nm. Calculate the energy of 2 photons of red light. (7 pts)
Answer #1

1) 0.025 Kg/L

1 Kg = 10^5 centigrams

1 L = 10 deciliter

So, 0.025 Kg/L = 0.025*10^5/10 = 250 cg/dL ... Answer

2) Volume of ethanol = Mass/density = 5.7/0.997 = 5.7172 ml

1 ml = 1000 mm^3

So, 5.7172 ml = 5.7172 * 1000 = 5717.15 mm^3 .... Answer

3)

P1 = 1.54 atm, V1 = 0.650 ml , T1 = 12+273 = 285 K

P2 = 1.01 atm , V2 = ? , T2 = 16+273 = 289 K

Now using ideal gas equation,

P1 * V1 / T1 = P2 * V2 / T2

V2 = P1 * V1 * T2 / T1 * P2

V2 = 1.54 * 0.650 * 289 / 285 * 1.01

V2 = 1.005 ml OR 1.0 ml ...Answer

4) We know that, energy of a photon is given by,

E = h * c / L

h is planck constant, c is speed of light and L is wavelength in meters.

E = (6.6*10^-34) * (3*10^8) / (650*10^-9)

E = 0.030462 * 10^-17

E = 3.0462 * 10^-19 J

Energy of 2 photons = 2 * E = 2 * 3.0462*10^-19 = 6.0924*10^-19 J .... Answer

Let me know in case of any concerns.